# Do Not (Always) Look Right: Investigating the Capabilities of Decoder-Based Large Language Models for Sequence Labeling

### 25 Jan 2024 v1

@@ Abstract
Pre-trained language models based on masked language modeling (MLM) objective excel in natural language understanding (NLU) tasks. While fine-tuned MLM-based encoders consistently outperform causal language modeling decoders of comparable size, a recent trend of scaling decoder models to multiple billion parameters resulted in large language models (LLMs), making them competitive with MLM-based encoders. Although scale amplifies their prowess in NLU tasks, LLMs fall short of SOTA results in information extraction (IE) tasks, many framed as sequence labeling (SL). However, whether this is an intrinsic limitation of LLMs or whether their SL performance can be improved remains unclear. To address this, we explore strategies to enhance the SL performance of “open” LLMs (Llama2 and Mistral) on IE tasks. We investigate bidirectional information flow within groups of decoder blocks, applying layer-wise removal or enforcement of the causal mask (CM) during LLM fine-tuning. This approach yields performance gains competitive with SOTA SL models, matching or outperforming the results of CM removal from all blocks. Our findings hold for diverse SL tasks, proving that “open” LLMs with layer-dependent CM removal outperform strong MLM-based encoders and instruction-tuned LLMs. However, we observe no effect from CM removal on a small scale when maintaining an equivalent model size, pre-training steps, and pre-training and fine-tuning data.
Do Not (Always) Look Right: Investigating the Capabilities of Decoder-Based Large Language Models for Sequence Labeling
@@ 1 Introduction
Pre-trained language models (PLMs) built upon the transformer architecture have demonstrated exceptional performance across many natural language understanding (NLU) tasks (Radford et al., 2018; Devlin et al., 2019; Yang et al., 2019; Clark et al., 2020; Raffel et al., 2020). Typically, achieving SOTA results in tasks such as sequence classification and sequence labeling involved a two-step process: pre-training on unlabeled corpora, followed by fine-tuning on task-specific data – a process often referred to as transfer learning (Raffel et al., 2020). Two prevailing architectures emerged, each coupled with a compatible pre-training paradigm: (1) the decoder architecture, utilizing causal language modeling (CLM) as its pre-training objective, and (2) the encoder architecture, with masked language modeling (MLM) pre-training objective.
@@
In direct comparisons with a similar number of model parameters and employing a transfer learning strategy, MLM-based encoders consistently outperformed CLM-based decoders on NLU tasks (Devlin et al., 2019). However, a shift occurred within the NLP community when encoder models ceased being scaled to the same magnitude of parameters and pre-training data as their decoder counterparts. Instead, there has been a pronounced trend toward scaling decoder models to multiple billion parameters, which led to large language models (LLMs). LLMs boosted the performance of many NLU tasks by combining their text generation capabilities with prompting techniques, eliminating the need for fine-tuning model parameters (Liu et al., 2023).
@@
Nevertheless, there is still considerable room for improvement, even for the largest decoder models, such as ChatGPT, which falls far behind SOTA results on fundamental NLU tasks. This holds particularly for information extraction (IE) tasks (Han et al., 2023), such as named entity recognition (NER), aspect-based sentiment analysis (ABSA), and event extraction (EE). These tasks are often cast as sequence labeling (token classification) tasks, and tackling sequence labeling (SL) by prompting LLMs has been proven quite difficult (Wang et al., 2023a). However, it is unclear whether the subpar performance of LLMs on SL tasks is their innate limitation or if it can be remedied. As the community stopped scaling up encoders, LLMs are solidifying their position as the de facto standard in the field. Thus, improving the SL performance of LLMs stands as the sole viable option.
@@
While an investigation of LLM capabilities for SL has practical merit, the challenge is that many LLMs are locked behind paywalls, and their weights are not publicly accessible. Fortunately, the community has been training and publicly releasing LLMs with multiple billions of parameters, such as Llama2 (Touvron et al., 2023), Mistral (Jiang et al., 2023), and OPT (Zhang et al., 2022). Leveraging parameter-efficient fine-tuning techniques and quantization, including QLoRA (Dettmers et al., 2023), facilitates experimentation with fine-tuning these models on task-specific data.These strategies allow leveraging “open” LLM’s hidden states for task-specific classifiers, enabling fine-tuning on consumer-grade hardware with comparable or better performance than full fine-tuning.
@@
Using the hidden state of the decoder’s last token serves as a reliable feature for classifying the input sequence. However, for SL, the attempts to utilize each token’s hidden state as a feature do not yield satisfying results. The causal mask, integral to endowing decoders with text generation capabilities, limits bidirectional information flow, preventing the model from “looking right,” i.e., attending to tokens in the sequence positioned to the right of the current token. For many SL tasks, the token’s label depends on the succeeding tokens, and omitting this context often results in subpar performance. Recently, Li et al. (2023) found that completely removing the causal mask from the Llama2 model during fine-tuning substantially enhances SL performance on NER.
@@
In this paper, we explore strategies to enhance the SL performance of “open” LLMs on IE tasks. Building on insight from Li et al. (2023), we explore bidirectional information flow within different decoder blocks. We find that layer-wise causal mask removal across LLM blocks can result in gains competitive with SOTA models, contrasting the approaches of removing or keeping the causal mask in all layers. Our analysis on a series of IE tasks (NER, ABSA, and EE) shows that layer-wise “looking right” is highly task-dependent. We compare against strong encoder-based sequence taggers and instruction-tuned LLMs. “Open” LLMs without causal masks in specific layers outperform these baselines in almost all scenarios. Finally, we draw comparisons with encoder-based models, assessing the relative strengths of the decoder and encoder architectures regarding parameter scale, training data, and the particular SL task. To rigorously evaluate the architectures, we pre-train small encoder and decoder models from scratch with an identical number of parameters on the same data. Our findings indicate the superiority of the same scale encoder architecture over the decoder on SL tasks, which is consistent even when the causal mask is removed while fine-tuning the small pre-trained decoder models. This suggests that the observed SL performance boost in LLMs upon removing the causal mask stems from scaling up.
@@ 2 Background and Related Work
The first decoder-based PLM with CLM pre-training that gained traction was GPT (Radford et al., 2018), while the first encoder-based PLM, with grand MLM pre-training success, was BERT (Devlin et al., 2019). Pre-training with CLM teaches the model to generate text and does not allow it to use the sequence context right to the token being processed. This is enabled via a causal mask. While beneficial and necessary for coherent text generation, restricting the model to look right is detrimental if we want to use the decoder to produce token-level embeddings. Encoders pre-trained with MLM and without the causal mask are much more successful in solving SL tasks because of the enabled bidirectional information flow. Although decoder-based models are bad at labeling tokens in a sequence, they can be used as successful sequence classifiers where the representation of the last token in the sequence is standardly leveraged for classification. Many decoder models implement sequence classification heads on top of decoder blocks, but not SL heads (Wolf et al., 2020). Li et al. (2023) found that removing the causal mask from all decoder layers of Llama2 (Touvron et al., 2023) improves the performance across sequence classification tasks and the IE task of named entity recognition (NER). However, Li et al. (2023) do not examine in depth why this phenomenon occurs, whether it occurs with layer-wise causal mask removal, and whether it is limited to the NER task. We build on this and inspect the effect of different unmasking configurations across layers of models with Llama decoder blocks, yielding interesting findings for IE and SL tasks. Furthermore, we train smaller decoders with CLM and encoders with MLM to demonstrate that the superiority of unmasked Llama2 is a consequence of the model and training data scale. The authors of Megatron-LM (Shoeybi et al., 2019) showed that scaling the BERT model to four billion parameters with MLM and the GPT-2 model to eight billion parameters with CLM brings gains on NLU tasks. Nevertheless, the community stopped scaling encoders and continued scaling decoders. This shift in strategy has given rise to notable properties in decoder-based LLMs, including enhanced zero- and few-shot learning capabilities (Wei et al., 2021; Brown et al., 2020), the ability for in-context learning (ICL) (Brown et al., 2020; Liu et al., 2023), and chain-of-thought prompting (Wei et al., 2022). Examples of LLMs with these emerging properties are GPT-3 (Brown et al., 2020), PaLM (Chowdhery et al., 2023), and ChatGPT.
@@
Many IE tasks, such as NER, subtasks of ABSA, and EE, can effectively be treated as SL tasks. Traditionally, these tasks were addressed with encoder models, mapping each token’s hidden state to task label space (Zhang et al., 2020; Zhou and Chen, 2021; Pontiki et al., 2015). However, this approach necessitates training a separate model for each task. More recent methods aim to infer the underlying structure of multiple IE tasks simultaneously and develop universal information extraction (UIE) systems, employing encoder, decoder, or encoder-decoder architectures. Often, the models of choice for this approach are encoder-decoder PLMs such as BART (Lewis et al., 2020) and T5 (Raffel et al., 2020). Building on T5 and instead of training task-specific classifiers, Paolini et al. (2021) propose the TANL framework, treating IE tasks as translation tasks between augmented natural languages. Similarly, Lu et al. (2022) introduce UIE, a text-to-structure generation framework leveraging T5. The framework generates target extractions through a schema-based prompt mechanism, with the structured schema learned in an unsupervised fashion. Further, the authors of LasUIE (Fei et al., 2022) leverage syntactic structure information for UIE improvement. Meanwhile, Wang et al. (2022a) extend the GLM (Du et al., 2022) with structure pre-training, proposing DeepStruct and reformulating structure prediction as a series of triple prediction tasks. Recently, Ping et al. (2023) revealed UniEX, a paradigm for UIE, utilizing BERT-like encoders to convert text-based IE into a token-pair problem. This is achieved by modifying the attention mechanism to extract spans in a unified manner. Another encoder-based approach recasts IE tasks into a multi-span cyclic graph extraction problem to end up with a non-autoregressive graph decoding algorithm, proving beneficial for zero and few-shot settings (Zhu et al., 2023). Finally, Lou et al. (2023) reconceptualize IE as a triplet prediction problem through semantic matching for UIE. We argue that the benefits of having a unified IE framework are undeniable. However, including more IE tasks or the same tasks with domain-specific data into each framework would still require re-training the UIE extractors. Therefore, we stick to the one model per IE and SL task approach, relying on parameter-efficient methods that yield compact and extensible models (Houlsby et al., 2019).
@@
IE tasks can be tackled with ICL, relying on the emergent properties of LLMs. Nevertheless, Pang et al. (2023) report that the performance of ICL lags behind SOTA supervised expert models. They develop a model that learns to guide the prompt so that LLMs output better results. Another example is GPT-NER (Wang et al., 2023a), where the authors transform the NER task into a generation one utilizing ICL and ChatGPT.
@@
Releasing the weights of encoder-decoder and decoder-only pre-trained LLMs paved the way for instruction tuning (IT) (Mishra et al., 2022). This approach involves fine-tuning the model to address specific tasks in a supervised generative manner, providing the model with instructions for each data point. The authors of InstructionNER (Wang et al., 2022b) propose a framework for low-resource NER. They reformulate the task as a generation problem with better prompt templates and instruction-tune the T5 model, stating that current prompt templates are primarily designed for sentence-level tasks and inappropriate for SL objectives. Extending the fusion of UIE and IT, InstructUIE (Wang et al., 2023b) reveals that multi-task IT can give results comparable to BERT (Devlin et al., 2019) in a supervised setting. Finally, Scaria et al. (2023) find that instruction-tuned encoder-decoder PLMs excel at ABSA subtasks. They used the Tk-instruct model, which is the PLM T5 fine-tuned in a supervised manner on various instruction tasks (Wang et al., 2022c).
@@ 3 Layer Group Unmasking
The causal mask is crucial in decoders, preventing the model from attending to future tokens and facilitating autoregressive text generation. This constraint is implemented by adding the causal triangular mask, denoted as CM, to the dot product of the query and key attention matrices. The resulting sum is passed through the softmax function in the scaled dot-product attention mechanism, as introduced by Vaswani et al. (2017). Formally, CM and the scaled dot-product attention are given as follows:
@@
Here, $Q$, $K$, and $V$ represent the query, key, and value attention matrices, respectively, and $d_{k}$ is the dimension of queries and keys. Effectively, applying softmax for the tokens with value $\infty$- ∞ in the CM results in attention scores of 00. To allow the LLM to “sometimes look right” (hence the title of our paper), we strategically choose in which decoder blocks to replace all $\infty$- ∞ entries in the CM with 00. This equates to the CM removal (actually, we leave the positions of padding tokens to $\infty$- ∞, but replace all others with 00). As Li et al. (2023) observed, removing CM across all decoder layers of Llama2 during fine-tuning increases the model’s performance on the NER SL task by a large margin. This is surprising because the model was pre-trained with the restriction to “look right” and can learn to attend to future tokens by backpropagating on the training data over a few epochs when the CM is removed. Motivated by this finding, we try to boost the performance of “open” LLMs further, trying out combinations of bidirectional information flow enablement (“unmasking”) or restriction (“masking”) per layer groups consisting of $b$ decoder blocks. We refer to these decisions as layer group “unmasking configurations.” Considering that Llama2 is made up of $n=32$ decoder blocks, and each block can either remove or enforce CM, we deal with a search space of $2^{n}=2^{32}$ possibilities. To limit the search space, we group the $n$ decoder blocks into $m=4$ layer groups with $b=8$ decoder blocks per layer group. Here, we only consider unmasking or masking all blocks per layer group. This leaves us with $162^{m}=$ 16 configurations per LLM and task. We encode our unmasking configurations with binary four-digit codes ranging from $0000$ (all masked) to $1111$ (all unmasked) where each 00 in the code represents a layer group with $b$ masked blocks, and vice-versa, each $1$ represents a layer group with $b$ unmasked blocks. These unmasking configurations are interpreted left-to-right, where the first digit in the code refers to the first $b$ decoder blocks closest to the model’s input.

### 27 Feb 2024 v2

# Looking Right is Sometimes Right: Investigating the Capabilities of Decoder-only LLMs for Sequence Labeling

@@ Abstract
Pre-trained language models based on masked language modeling (MLM) excel in natural language understanding (NLU) tasks. While fine-tuned MLM-based encoders consistently outperform causal language modeling decoders of comparable size, recent decoder-only large language models (LLMs) perform on par with smaller MLM-based encoders. Although their performance improves with scale, LLMs fall short of achieving state-of-the-art results in information extraction (IE) tasks, many of which are formulated as sequence labeling (SL). We hypothesize that LLMs’ poor SL performance stems from causal masking, which prevents the model from attending to tokens on the right of the current token. Yet, how exactly and to what extent LLMs’ performance on SL can be improved remains unclear. We explore techniques for improving the SL performance of open LLMs on IE tasks by applying layer-wise removal of the causal mask (CM) during LLM fine-tuning. This approach yields performance gains competitive with state-of-the-art SL models, matching or outperforming the results of CM removal from all blocks. Our findings hold for diverse SL tasks, demonstrating that open LLMs with layer-dependent CM removal outperform strong MLM-based encoders and even instruction-tuned LLMs.
Looking Right is Sometimes Right: Investigating the Capabilities of Decoder-only LLMs for Sequence Labeling
@@ 1 Introduction
Pre-trained language models (PLMs) built upon the Transformer architecture have demonstrated exceptional performance across many natural language understanding (NLU) tasks (Radford et al., 2018; Devlin et al., 2019; Yang et al., 2019; Clark et al., 2020; Raffel et al., 2020). Typically, achieving state-of-the-art (SOTA) results in tasks such as sequence classification and sequence labeling involves a two-step process: pre-training on unlabeled corpora, followed by fine-tuning on task-specific data – a process often referred to as transfer learning (Ruder et al., 2019; Raffel et al., 2020). Two prevailing architectures emerged, each coupled with a compatible pre-training paradigm: (1) the decoder-only architecture, utilizing causal language modeling (CLM) for pre-training, and (2) the encoder-only architecture, with the masked language modeling (MLM) pre-training objective.
@@
In transfer learning experiments that juxtapose models of a comparable number of parameters, MLM-based encoders consistently outperformed CLM-based decoders on NLU tasks (Devlin et al., 2019). [^1] Henceforth, we use the terms “encoder” and “decoder” to refer exclusively to MLM- and CLM-based encoder- and decoder-only variants, respectively. However, a shift in strategy emerged within the NLP community when encoder models ceased being scaled up to the same magnitude of parameters and pre-training data as their decoder counterparts. Consequently, there has been a pronounced trend toward scaling decoder models to multiple billion parameters, leading to a proliferation of large language models (LLMs). Combining LLM text generation capabilities with various prompting strategies can boost the performance on many NLU tasks, eliminating the need for fine-tuning model parameters (Liu et al., 2023).
@@
In spite of LLMs’ good performance on NLU tasks, there is still considerable room for improvement, even for the largest decoder models, such as ChatGPT, which fall far behind SOTA results on fundamental NLP tasks. This holds particularly for information extraction (IE) tasks (Han et al., 2023), such as named entity recognition (NER), aspect-based sentiment analysis (ABSA), and event extraction (EE). These tasks are often formulated as sequence labeling (SL) tasks, and tackling SL by prompting LLMs proved quite difficult (Wang et al., 2023a). However, it is unclear how exactly and to what extent the subpar performance of LLMs on SL tasks can be remedied. As the community stopped scaling up encoders, LLMs are solidifying their position as field’s de facto standard. Thus, improving LLMs’ SL performance is the sole viable option.
@@
Although many SOTA LLMs are accessible only through paywalls, the community has responded by training and publicly releasing open LLMs with multiple billions of parameters, such as Llama2 (Touvron et al., 2023), Mistral (Jiang et al., 2023), and OPT (Zhang et al., 2022). Parameter-efficient fine-tuning (PEFT) techniques and quantization, including QLoRA (Dettmers et al., 2023), facilitate experimentation with fine-tuning these models on task-specific data. These strategies allow leveraging open LLMs’ hidden states for task-specific classifiers, enabling fine-tuning on consumer-grade hardware with comparable or better performance than full fine-tuning.
@@
The poor performance of LLMs on SL tasks can be traced back to the information encoded in the hidden states. Using the hidden state of the decoder’s last token serves as a reliable feature for classifying the input sequence. However, the same tactic does not suffice for SL because the causal mask (CM), integral to endowing decoders with text generation capabilities, limits bidirectional information flow, preventing the model from “looking right,” i.e., attending to tokens in the sequence positioned to the right of the current token. For many SL tasks, the token’s label depends on the succeeding tokens, and omitting this context often results in subpar performance. Recently, Li et al. (2023) found that completely removing the CM from the Llama2 model during fine-tuning substantially enhances SL performance on NER.
@@
In this paper, we explore techniques to enhance the SL performance of open LLMs on IE tasks. Building on insight from Li et al. (2023), we hypothesize that removing the CM from all decoder layers may not be beneficial for all SL tasks. To confirm this, we experiment with layer-wise CM removal across LLM blocks, observing gains competitive with SOTA models, in contrast to removing or keeping the CM in all layers. We extend our analysis to a series of IE tasks (NER, ABSA, and EE), and demonstrate that layer-wise “looking right” is highly task-dependent. We compare against strong encoder-only sequence taggers and instruction-tuned LLMs. Open LLMs without CMs in specific layers outperform these baselines in almost all scenarios. Finally, we compare with encoder-only models, assessing the relative strengths of the decoder and encoder architectures regarding parameter scale, training data, and the particular SL task. We pre-train small encoder and decoder models from scratch with an identical number of parameters on the same data. Our findings indicate the superiority of the same-scale encoder architecture over the decoder on SL tasks, which is consistent even when the CM is removed while fine-tuning the small pre-trained decoders. This suggests that the observed SL performance gain in LLMs upon CM removal stems from scaling up.
Our contributions are threefold: (1) We present evidence that layer-wise CM removal, followed by supervised fine-tuning, improves the performance of decoder-only LLMs on SL tasks compared to removing or keeping CM in all layers; (2) Layer-wise CM removal configuration that yields the highest gains strongly depends on the peculiarities of the SL task; (3) We show that the CM removal effect brings no gains on the small scale. We believe our results will contribute to building more performant decoder-only SL models.
@@ 2 Background and Related Work
The first encoder-only PLM, with MLM pre-training success, was BERT (Devlin et al., 2019), while the first decoder-only PLM with CLM pre-training that gained traction was GPT (Radford et al., 2018). Shoeybi et al. (2019) showed that scaling the BERT model to four billion parameters with MLM and the GPT-2 model (Radford et al., 2019) to eight billion parameters with CLM brings gains on NLU tasks. Nevertheless, the community stopped scaling up encoders and continued scaling decoders. This shift in strategy has given rise to notable properties in decoder-only LLMs, including enhanced zero- and few-shot learning capabilities (Wei et al., 2021; Brown et al., 2020), the ability for in-context learning (ICL) (Brown et al., 2020; Liu et al., 2023), and chain-of-thought prompting (Wei et al., 2022). Examples of LLMs with these emerging properties are GPT-3 (Brown et al., 2020), PaLM (Chowdhery et al., 2023), and ChatGPT.
Pre-training with CLM teaches the model to generate text and does not allow it to use the sequence context right to the token being processed. This is enabled via a CM, which prevents the model from attending to future tokens. While beneficial and necessary for coherent text generation, restricting the model to look right is detrimental if one wants to use the decoder to produce fully contextualized token-level embeddings with the decoder. Although decoder-only models perform poorly on sequence labeling tasks (i.e., tasks where each individual token is assigned a label), they perform quite well on sequence classification tasks (i.e., tasks where the entire sequence is assigned one label). Therefore, many decoder models implement sequence classification heads on top of decoder blocks, but not SL heads (Wolf et al., 2020).
Li et al. (2023) found that removing the CM from all decoder layers of Llama2 (Touvron et al., 2023) improves the performance of named entity recognition (NER). However, Li et al. (2023) did not examine in depth why this phenomenon occurs, whether it occurs with layer-wise CM removal, and whether it persists for other IE tasks. We build on this and inspect the effect of different unmasking configurations across layers of open decoder-only LLMs. Furthermore, we train small decoders with CLM and small encoders with MLM to demonstrate that the superiority of unmasked LLMs is a consequence of the model and training data scale.
@@
Many IE tasks, such as NER, subtasks of ABSA, and EE, can effectively be formulated as SL tasks. Traditionally, these tasks were addressed with encoders, mapping each token’s hidden state to task label space with a linear layer on top of the last encoder block (Pontiki et al., 2015; Zhang et al., 2020; Zhou and Chen, 2021). However, this approach necessitates training a separate model for each task. More recent methods aim to infer the underlying structure of multiple IE tasks simultaneously and develop universal information extraction (UIE) models, employing encoder, decoder, or encoder-decoder architectures (Paolini et al., 2021; Lu et al., 2022; Fei et al., 2022; Wang et al., 2022a; Ping et al., 2023; Zhu et al., 2023; Lou et al., 2023). While the practical advantages of UIE are obvious, including more IE tasks would still require re-training the UIE extractors. Therefore, we stick to the one model per SL task strategy, relying on parameter-efficient methods that yield compact and extensible models (Houlsby et al., 2019).
@@
ICL is a powerful method for scrutinizing LLMs, employed by prompting LLMs with task demonstrations. IE tasks can be tackled with ICL, relying on the emergent properties of LLMs. Nevertheless, Pang et al. (2023) report that the performance of LLMs leveraging ICL lags behind the SOTA results of supervised IE models. They develop a model that learns to guide the prompt for improved LLM performance. Another notable example is GPT-NER (Wang et al., 2023a), transforming the NER task into a generation one utilizing ICL and ChatGPT.
@@
Releasing the weights of encoder-decoder and decoder-only pre-trained LLMs paved the way for instruction tuning (IT) (Mishra et al., 2022). This paradigm involves fine-tuning the model to address specific tasks in a supervised generative manner, providing the model with instructions for each data point. SL tasks can be adapted for compatibility with IT. Arguing that current prompt templates are primarily designed for sentence-level tasks and inappropriate for SL objectives, Wang et al. (2022b) reformulate NER as a generation problem relying on improved prompt templates and IT. Drawing on the fusion of UIE and IT, Wang et al. (2023b) reveal that multi-task IT can give results comparable to BERT in a supervised setting for NER. Finally, Scaria et al. (2023) find that instruction-tuned encoder-decoder PLMs excel at ABSA subtasks, leveraging the Tk-instruct model (Wang et al., 2022c). Given that IT is a powerful supervised paradigm for combining LLMs’ generation abilities with labeled data, we compare SL models against IT baselines.
@@ 3 Layer Group Unmasking
The CM is a crucial component of CLM-based decoders, preventing the model from attending to future tokens and facilitating autoregressive text generation. This constraint is enforced by defining the CM as a triangular matrix and adding this matrix to the dot product of the query and key attention matrices. The resulting sum is passed through the softmax function in the scaled dot-product attention mechanism, as introduced by Vaswani et al. (2017). Formally:
@@
Here, $Q$, $K$, and $V$ are the query, key, and value attention matrices, respectively, and $d_{k}$ is the dimension of queries and keys. Effectively, applying softmax for tokens with value $\infty$- ∞ in the CM results in attention scores being 0.
As reported by Li et al. (2023), removing CM across all decoder layers of Llama2 during fine-tuning increases the model’s performance on the NER SL task by a large margin. This is surprising because although, during pre-training, the CM mask was in place and the model was restricted from “looking right,” it learned to attend to future tokens by backpropagating on the training data over a few epochs with CM removed. Building on this and to allow the LLM to “sometimes look right,” we select a subset of decoder blocks for which we replace all $\infty$- ∞ entries in the CM with zeros, effectively removing the CM (technically, we leave the positions of padding tokens to $\infty$- ∞, but replace all others with 0). Since we group layers into groups of $b$ blocks, we refer to these decisions as layer group unmasking configurations.
Considering that Llama2 is made up of $n=32$ decoder blocks, and each block can either remove or keep its CM, this gives a search space of $2^{n}=2^{32}$ possibilities. To reduce the search space to a manageable size, we group the $n$ decoder blocks into $m=4$ layer groups with $b=8$ consecutive decoder blocks per layer group and then jointly mask or unmask all blocks in each layer group. This leaves us with $162^{m}=$ 16 unmasking configurations per LLM and task. For ease of reference, we encode our unmasking configurations with binary four-digit codes ranging from $0$ (all four layer groups masked) to $1$ (all four layer groups unmasked), where each 0 and 1 denote a layer group that is masked or unmasked, respectively. Unmasking configurations are interpreted left to right (the first digit pertains to the layer group closest to the model’s input).